OpenStudy (anonymous):

Your cap sixe is based on the circumference of your head (in inches). For a certain range of head circumferences c, cap size s can be modeled by the following equation. S= c-1/3 If you wear a size 6.5 cap, what is your head circumference?

OpenStudy (anonymous):

Size= (circumference -1)/3 7=(c-1)/3 c-1=21 c=22 cm